two identical 15 kg balls each 25 cm on diameter are suspended by two 35 cm wires. The entire apparatus is suspended by a single 18 cm wire. Find the tension in each of the 3 wires


1
Expert's answer
2021-03-04T11:55:58-0500

"T_1=T_2=T"


"T\sin\alpha=mg\to T=\frac{mg}{\sin\alpha}"


"\sin \alpha=\frac{\sqrt{0.35^2-0.125^2}}{0.35}=0.934"


"T=\frac{mg}{\sin\alpha}=\frac{15\cdot9.8}{0.934}\approx157\ (N)" . Answer


"T_3=2mg=2\cdot15\cdot9.8=294\ (N)" . Answer